# VisNumBench: Evaluating Number Sense of Multimodal Large Language Models

### v1

@@ Abstract
Can Multimodal Large Language Models (MLLMs) develop an intuitive number sense similar to humans? Targeting this problem, we introduce Visual Number Benchmark (VisNumBench) to evaluate the number sense abilities of MLLMs across a wide range of visual numerical tasks. VisNumBench consists of about multiple-choice question-answer pairs derived from both synthetic and real-world visual data, covering seven visual numerical attributes and four types of visual numerical estimation tasks. Our experiments on VisNumBench led to the following key findings: (i) The 17 MLLMs we tested—including open-source models such as Qwen2.5-VL and InternVL2.5, as well as proprietary models like GPT-4o and Gemini 2.0 Flash—perform significantly below human levels in number sense-related tasks. (ii) Multimodal mathematical models and multimodal chain-of-thought (CoT) models did not exhibit significant improvements in number sense abilities. (iii) Stronger MLLMs with larger parameter sizes and broader general abilities demonstrate modest gains in number sense abilities. We believe VisNumBench will serve as a valuable resource for the research community, encouraging further advancements in enhancing MLLMs’ number sense abilities. All benchmark resources, including code and datasets, will be publicly available at https://wwwtttjjj.github.io/VisNumBench/.
@@ 1 Introduction
VNS-Bench | Angle | Length | Scale | Quantity | Depth | Area | Volume | Total |
VNS-Bench-Synthetic | 170 | 181 | 140 | 196 | 135 | 189 | - | 1011 |
VNS-Bench-Real | 149 | 162 | 143 | 147 | 154 | - | 147 | 902 |
@@
In this work, we introduce the Visual Number Benchmark (VisNumBench), inspired by human number sense abilities. As illustrated in Figure 2, VisNumBench is structured into two components based on different visual scenarios: VisNumBench-Synthetic and VisNumBench-Real. VisNumBench-Synthetic comprises controlled, synthetic images in which numerical relationships are explicitly defined. VisNumBench-Real contains real-world images, providing a more complex and less controlled environment. Grounded in human number sense, VisNumBench targets seven key dimensions of visual numerical attributes through four distinct types of visual numerical estimation tasks, as depicted in Figure 3.
We evaluated 17 MLLMs on VisNumBench and found that even state-of-the-art models perform poorly on our proposed benchmark, which is shown in Figure 4. Furthermore, our experiments reveal that adopting multimodal mathematical models and multimodal Chain-of-Thought (CoT) models did not lead to substantial performance improvements. However, the performance of the latest models is better than the previous models from the same family. For example, Qwen2.5VL [41] performs better than Qwen2VL [44]. It seems that optimization on data, training techniques, and model architecture will help models improve number sense ability. Within this work, we aim at advancing MLLMs toward higher levels of intelligence by developing models that enhance visual number sense abilities.
The main contributions of this paper are listed as follows:
@@
We introduce VisNumBench, a comprehensive benchmark that integrates diverse data sources and an automated evaluation framework to assess the numerical sense abilities of MLLMs across various visual numeric tasks.
@@
We conduct a comprehensive evaluation of various MLLMs on VisNumBench and find that even the most advanced models still demonstrate limited numerical sense abilities.
@@ 2.2 Benchmarks for MLLMs
Advancements in MLLMs have led to the development of numerous benchmarks aimed at evaluating model performance across a broad spectrum of general multimodal tasks. Several recent studies [15, 53, 28, 22, 21, 52] have introduced more comprehensive multimodal benchmarks that provide extensive and holistic assessments. Beyond general multimodal tasks, specialized benchmarks have been created to evaluate the mathematical reasoning capabilities of MLLMs. Tasks such as visual reasoning with numbers, arithmetic problem-solving, and algebraic manipulation play a crucial role in assessing both the numerical proficiency and higher-order cognitive abilities of MLLMs. Prominent benchmarks, including MATHVISTA [29], Math-Vision [43], MathOdyssey [13], and SMART-840 [9], are designed to test models on a diverse range of mathematical challenges, such as word problems, equation-solving, and complex multi-step reasoning.
@@ 3.2.1 Source of Images
Web Collection. To incorporate more natural and diverse visual data, we gathered figures containing numerical information from publicly available web resources [42, 17, 12]. These figures were carefully curated and filtered to ensure relevance and clarity before designing the corresponding questions.
@@
3.2.2 Construction of QA Pairs and Data Quality Control
As shown in parts (b) and (c) in Figure 5, we involve QA pairs for images from different sources. For the figures generated by the Python program, we manually design different questions based on the specific characteristics of each figure and generate corresponding annotations using the parameters saved We can design the question such as: “In the ABCD sections of the pie chart, which section has the largest/smallest area?” Based on the parameters saved, the answer would be “C / D”. For images from other sources, we manually designed questions and annotated them based on different numerical attributes of the images. In addition, we can generate different numerical estimation tasks based on the answer type. For example, for the question: “Which of the following options is a reasonable estimate of the number of cubes in the figure?”, when the answer is “(50, 75)”, it corresponds to a range estimation task. On the contrary, if the answer is “62”, it belongs to the value estimation task.
We employ a combination of automated and manual methods to design distractors. For numerical answers, we generate alternative options that are easily confusable with the correct answer. Some distractors are constructed based on their inherent properties (e.g., areas A, B, and D in part (c) of Figure 5). These distractors are selected to align with human perceptual biases, ensuring that they appear plausible while remaining distinguishable from the correct answer.
@@ 4 Experiments
We randomly selected samples ( QA pairs from each category), with sourced from VisNumBench-Synthetic and from VisNumBench-Real. Human evaluators independently answered each question and provided assessments. Accuracy (%) is reported for all experimental results, and all the results are provided in Tables 2 and 3.
@@ 4.1 Evaluation Results and Analysis
From Tables 2 and 3, we observe that the performance of MLLMs is not comparable to that of humans. Among the seven types of questions, quantity-related tasks appear to be the easiest, while angle-related tasks are the most difficult. This is possibly because the amount of data available for quantity-related tasks is significantly greater than that for angle-related tasks. By comparing the evaluations on synthetic and real images, we find that the performance of the same model does not exhibit significant variance. Thus, in terms of numerical reasoning ability, both synthetic and real images present similar challenges for existing MLLMs. Moreover, we observe that the best open-sourced model performs comparably to the best closed-source models. More detailed analyses and discussions are provided in the following subsections.
@@ 4.1.2 Performance on VisNumBench-Real
Accuracy on the VisNumBench-Real dataset shows similar trends. InternVL2.5-78B and Gemini 2.0 Flash stand out with an average accuracy of , achieving near-optimal results across multiple tasks, as shown in Table 3. InternVL2.5-38B attained an exceptionally high accuracy of on the quantity task, while LLaVA-v1.6-34B excelled in the volume task, achieving the highest scores. Qwen2.5-VL-72B demonstrated relatively balanced performance across tasks, yielding a suboptimal average accuracy of .
@@ 4.1.3 Performance on Different Visual Numerical Estimation Tasks
As we analyze performance across different numerical estimation tasks, Figure 6 reveals that in the synthetic scenario, Gemini 2.0 Flash and Qwen2.5-VL-72B achieve the highest performance across all sub-tasks, particularly in range estimation, value estimation, and value comparison, where their scores consistently exceed 60. In contrast, GPT-4o exhibits the lowest performance in all tasks, especially in value comparison and multiplicative estimation. In the real-world scenario, most models achieved their best performance in value comparison tasks, which are also the easiest for humans. Although Gemini 2.0 Flash and InternVL2.5-78B continue to perform well in most tasks, their performance in multiplicative estimation has declined compared to the synthetic scenario. Additionally, GPT-4o continues to perform poorly across all tasks, particularly in multiplicative estimation and value comparison, where it falls significantly behind other models.
Notably, in the multiplicative estimation task, LLaVA-v1.6-34B outperforms all other models by a significant margin. This suggests that certain models may be more specialized for specific types of tasks, and further fine-tuning or adjustments could enhance performance across different tasks.
@@ 4.2 Further Analysis
How do math-special models perform on the VisNumBench? To investigate the number sense abilities of math-special models, we introduce two multimodal mathematical models: (1) InternVL2-8B-MPO [47], initialized from InternVL2-8B [8] and fine-tuned on the large-scale multimodal reasoning preference dataset MMPR [47], achieving the score of on MathVista; (2) Math-LLaVA-13B [39], initialized from LLaVA-v1.5-13B and fine-tuned on the MathV360K [39] dataset. As shown in Figure 7, InternVL2-8B-MPO achieved a improvement in synthetic scenarios and a increase in real-world scenarios. Its enhancements are task-specific rather than universally effective across different number sense challenges. In contrast, Math-LLaVA-13B exhibited a polarized performance trend: while it improved by on the synthetic dataset, its accuracy declined by in real-world scenarios. This suggests that although the model benefits from training on synthetic data, it struggles to generalize to the complexity and variability of real-world number sense tasks. Relying solely on synthetic data may be insufficient to enhance number sense capabilities in real-world applications. Additional strategies, such as incorporating more diverse real-world training data or refining model architectures, may be necessary to achieve meaningful improvements.
@@
How do the multimodal reasoning models perform? To examine whether reasoning techniques can enhance the number sense abilities of MLLMs, we evaluate two multimodal reasoning models: Llama-3.2V-11B-cot222https://huggingface.co/Xkev/Llama-3.2V-11B-cot [50] and R1-OneVision-7B333https://github.com/Fancy-MLLM/R1-Onevision. Llama-3.2V-11B-cot is trained using LLaVA-o1-100k [50], achieving a performance improvement on MathVista compared to Llama-3.2-11B-Vision-Instruct. R1-OneVision-7B, trained with a rule-based reinforcement learning technique, attains an accuracy of on Mathverse [54]. Accordingly, we assess these models on our benchmark. The results, presented in Figure 7, indicate that neither Llama-3.2V-11B-cot nor R1-OneVision-7B achieved the expected performance gains. On the contrary, their accuracy dropped significantly—except for a modest improvement by Llama-3.2V-11B-cot in synthetic scenarios—especially in real-world settings. These findings suggest that developing reasoning techniques specifically tailored for number sense abilities may be necessary.
@@ 6 Acknowledge
This research was supported by the National Natural Science Foundation of China No.62272315.
@@ Appendix A Appendix Outline
Appendix B Details for VisNumBench Construction
@@
Length Based on synthetic and real scenes, we designed various types of question-answering questions, such as relative length comparison and multiple estimation of line segments, as well as the estimation of length, height, and proportion of different objects, among others. The figures for VisNumBench-Synthetic are either generated by Python programs or sourced from MathVista [29], whereas the figures for VisNumBench-Real are captured by the authors or collected from Google Images [17].
@@
More details of data construction are shown in the Tables 5, 6. Figure 8 shows the dataset statistics of VisNumBench based on various visual numerical estimation tasks. Figure 9 shows how to build a QA pair based on images generated by a Python script.
@@ C.1 Model Access
This section details the model access and model parameters (refer to Table 7). The model responses presented in this paper were collected between January 1 and February 28, 2025.
@@
C.2 Additional Results
@@
Tablse 8 and 9 show the results by multimodal mathematical models and multimodal CoT models, corresponding to Figure 7. Table 10 presents the performance of models of varying sizes from the QwenVL and InternVL families, extending the results reported in Table 4. .
@@
C.3 Error Analysis
We take the results of Gemini 2.0 Flash as an example to conduct an error analysis. Based on these results, we randomly selected 10 erroneous instances for each attribute in each scenario and manually analyzed a total of 120 randomly sampled errors across all tasks. We categorize the errors into two types: (1) Errors resulting from the model’s failure to correctly perceive the image (image perception errors). (2) Errors where the model correctly interprets the image and the question but fails to generate the correct numerical answer (numerical intuition errors). Our findings indicate that 28.3% of the errors stem from image perception errors, particularly in scale-related tasks, where the model struggles to recognize the position of pointers. In 71.7% of cases, the model exhibits numerical intuition errors, which manifest in challenges related to depth estimation, angle relationships, and quantity perception. These findings further confirm that the current model indeed lacks robust numerical intuition.

### v2

@@ Abstract
Can Multimodal Large Language Models (MLLMs) develop an intuitive number sense similar to humans? Targeting this problem, we introduce Visual Number Benchmark (VisNumBench) to evaluate the number sense abilities of MLLMs across a wide range of visual numerical tasks. VisNumBench consists of about multiple-choice question-answer pairs derived from both synthetic and real-world visual data, covering seven visual numerical attributes and four types of visual numerical estimation tasks. Our experiments on VisNumBench led to the following key findings: (i) The 17 MLLMs we tested—including open-source models such as Qwen2.5-VL and InternVL2.5, as well as proprietary models like GPT-4o and Gemini 2.0 Flash—perform significantly below human levels in number sense-related tasks. (ii) Multimodal mathematical models and multimodal chain-of-thought (CoT) models did not exhibit significant improvements in number sense abilities. (iii) Stronger MLLMs with larger parameter sizes and broader general abilities demonstrate modest gains in number sense abilities. We believe VisNumBench will serve as a valuable resource for the research community, encouraging further advancements in enhancing MLLMs’ number sense abilities. Code and dataset are available at https://wwwtttjjj.github.io/VisNumBench/.
@@ 1 Introduction
| VisNumBench | Angle | Length | Scale | Quantity | Depth | Area | Volume | Total |
| VisNumBench-Synthetic | 170 | 181 | 140 | 196 | 135 | 189 | - | 1011 |
| VisNumBench-Real | 149 | 162 | 143 | 147 | 154 | - | 147 | 902 |
@@
In this work, we introduce the Visual Number Benchmark (VisNumBench), inspired by human number sense abilities. As illustrated in Figure 2, VisNumBench is structured into two components based on different visual scenarios: VisNumBench-Synthetic and VisNumBench-Real. VisNumBench-Synthetic comprises controlled, synthetic images in which numerical relationships are explicitly defined. VisNumBench-Real contains real-world images, providing a more complex and less controlled environment. VisNumBench targets seven key dimensions of visual numerical attributes through four distinct types of visual numerical estimation tasks, as depicted in Figure 3.
We evaluated 17 MLLMs on VisNumBench and found that even state-of-the-art models perform poorly on our proposed benchmark, which is shown in Figure 4. Furthermore, our experiments reveal that adopting multimodal mathematical models and multimodal Chain-of-Thought (CoT) models did not lead to substantial performance improvements. However, the performance of the latest models is better than the previous models from the same family. For example, Qwen2.5VL [42] performs better than Qwen2VL [45]. It seems that optimization on data, training techniques, and model architecture will help models improve their number sense ability. In this work, we aim to advance MLLMs toward higher levels of intelligence by developing models that enhance visual number sense abilities. The main contributions of this paper are listed as follows:
@@
We introduce VisNumBench, a comprehensive benchmark integrating diverse data sources and an automated evaluation framework to assess the numerical sense abilities of MLLMs across various visual numerical tasks.
@@
We conduct a comprehensive evaluation of various MLLMs on VisNumBench, finding that even the most advanced models still exhibit limited numerical sense.
@@ 2.2 Benchmarks for MLLMs
Advancements in MLLMs have led to the development of numerous benchmarks aimed at evaluating model performance across a broad spectrum of general multimodal tasks. Several recent studies [15, 54, 28, 22, 21, 53, 16, 20] have introduced more comprehensive reasoning and perception multimodal benchmarks that provide extensive and holistic assessments. Beyond general multimodal tasks, specialized benchmarks have been created to evaluate the mathematical reasoning capabilities of MLLMs. Tasks such as visual reasoning with numbers, arithmetic problem-solving, and algebraic manipulation play a crucial role in assessing both the numerical proficiency and higher-order cognitive abilities of MLLMs. Prominent benchmarks, including MATHVISTA [30], Math-Vision [44], MathOdyssey [13], and SMART-840 [9], are designed to test models on a diverse range of mathematical challenges, such as word problems, equation-solving, and complex multi-step reasoning.
@@ 3.2.1 Source of Images
3.2.2 QA Pair Construction and Quality Control
As shown in parts (b) and (c) in Figure 5, we involve QA pairs for images from different sources. For the figures generated by the Python program, we manually design different questions based on the specific characteristics of each figure and generate corresponding annotations using the parameters saved. We can design the question such as: “In the ABCD sections of the pie chart, which section has the largest/smallest area?” Based on the parameters saved, the answer would be “C / D”. For images from other sources, we manually designed questions and annotated them based on different numerical attributes of the images. In addition, we can generate different numerical estimation tasks based on the answer type. For example, for the question: “Which of the following options is a reasonable estimate of the number of cubes in the figure?”, when the answer is “(50, 75)”, it corresponds to a range estimation task. On the contrary, if the answer is “62”, it belongs to the value estimation task.
We employ a combination of automated and manual methods to design distractors. For numerical answers, we generate alternative options that are easily confusable with the correct answer. Some distractors are constructed based on their inherent properties (e.g., areas A, B, and D in part (c) of Figure 5). These distractors are designed to align with human perceptual biases, appearing plausible yet distinguishable from the correct answer.
@@ 4 Experiments
We randomly selected samples ( QA pairs from each numerical attribute), with sourced from VisNumBench-Synthetic and from VisNumBench-Real. Human evaluators independently answered each question and provided assessments. Accuracy (%) is reported for all experimental results, and all the results are provided in Tables 2 and 3.
@@ 4.1 Evaluation Results and Analysis
From Tables 2 and 3, we observe that the performance of MLLMs is not comparable to that of humans. Among the seven types of questions, quantity-related tasks appear to be the easiest, while angle-related tasks are the most difficult. This is possibly because the amount of training data available for quantity-related tasks is significantly greater than that for angle-related tasks. By comparing the evaluations on synthetic and real images, we find that the performance of the same model does not exhibit significant variance. Thus, in terms of numerical reasoning ability, both synthetic and real images present similar challenges for existing MLLMs. Moreover, we observe that the best open-sourced model performs comparably to the best closed-source models.
More detailed analyses and discussions are provided in the following subsections.
@@ 4.1.2 Performance on VisNumBench-Real
Accuracy on the VisNumBench-Real dataset shows similar trends. InternVL2.5-78B and Gemini 2.0 Flash stand out with an average accuracy of , achieving near-optimal results across multiple tasks, as shown in Table 3. InternVL2.5-38B attained an exceptionally high accuracy of on the quantity task, while LLaVA-v1.6-34B excelled in the volume task, achieving the highest scores. Qwen2.5-VL-72B demonstrated relatively balanced performance, yielding a suboptimal average accuracy of .
@@ 4.1.3 Performance on Different Visual Numerical Estimation Tasks
As we analyze performance across different numerical estimation tasks, Figure 6 reveals that in the synthetic scenario, Gemini 2.0 Flash and Qwen2.5-VL-72B achieve the highest performance across all visual numerical estimation tasks, particularly in range estimation, value estimation, and value comparison, where their accuracies consistently exceed 60%. In contrast, GPT-4o exhibits the lowest performance in all tasks, especially in value comparison and multiplicative estimation. In the real-world scenario, most models achieved their best performance in value comparison tasks, which are also the easiest for humans. Although Gemini 2.0 Flash and InternVL2.5-78B continue to perform well in most tasks, their performance in multiplicative estimation has declined compared to the synthetic scenario. Additionally, GPT-4o continues to perform poorly across all tasks, particularly in multiplicative estimation and value comparison, where it falls significantly behind other models.
Notably, in the multiplicative estimation task, LLaVA-v1.6-34B outperforms all other models by a significant margin. This suggests that certain models may be more specialized for specific types of tasks, and further fine-tuning could enhance performance across different tasks.
@@ 4.2 Further Analysis
How do math-special models perform on the VisNumBench? To investigate the number sense abilities of math-special models, we introduce two multimodal mathematical models: (1) InternVL2-8B-MPO [49], initialized from InternVL2-8B [8] and fine-tuned on the large-scale multimodal reasoning preference dataset MMPR [49], achieving an accuracy of 65.65% on MathVista; (2) Math-LLaVA-13B [40], initialized from LLaVA-v1.5-13B and fine-tuned on the MathV360K [40] dataset. As shown in Figure 7, InternVL2-8B-MPO achieved a improvement in synthetic scenarios and a increase in real-world scenarios. Its enhancements are task-specific rather than universally effective across different number sense challenges. In contrast, Math-LLaVA-13B exhibited a polarized performance trend: while it improved by on the synthetic dataset, its accuracy declined by in real-world scenarios. This suggests that although the model benefits from training on synthetic data, it struggles to generalize to the complexity and variability of real-world number sense tasks. Relying solely on synthetic data may be insufficient to enhance number sense capabilities in real-world applications. Additional strategies, such as incorporating more diverse real-world training data or refining model architectures, may be necessary to achieve meaningful improvements.
@@
How do the multimodal reasoning models perform? To examine whether reasoning techniques can enhance the number sense abilities of MLLMs, we evaluate two multimodal reasoning models: Llama-3.2V-11B-cot222https://huggingface.co/Xkev/Llama-3.2V-11B-cot [51] and R1-OneVision-7B333https://github.com/Fancy-MLLM/R1-Onevision. Llama-3.2V-11B-cot is trained using LLaVA-o1-100k [51], achieving a performance improvement on MathVista compared to Llama-3.2-11B-Vision-Instruct [34]. R1-OneVision-7B, trained with a rule-based reinforcement learning technique, attains an accuracy of on Mathverse [55]. Accordingly, we assess these models on our benchmark. The results, presented in Figure 7, indicate that neither Llama-3.2V-11B-cot nor R1-OneVision-7B achieved the expected performance gains. On the contrary, their accuracy dropped significantly—except for a modest improvement by Llama-3.2V-11B-cot in synthetic scenarios—especially in real-world settings. These findings suggest that developing reasoning techniques specifically tailored for number sense abilities may be necessary.
@@ 6 Acknowledge
This research was supported by the National Natural Science Foundation of China under Grant No. 62272315.
@@ Appendix A Appendix Outline
Appendix B Details of VisNumBench Construction
@@
Length Based on both synthetic and real-world scenes, we designed a variety of question-answering tasks, including relative length comparison, multiple segment estimation, and the estimation of object length, height, and proportion, among others. The figures in VisNumBench-Synthetic are either generated using Python programs or sourced from MathVista [30], whereas the figures in VisNumBench-Real are captured by the authors or collected from Google Images [17].
@@
More details of data construction are shown in the Tables 5, 6. Figure 8 shows the dataset statistics of VisNumBench based on various visual numerical estimation tasks. Figure 9 shows how to build a QA pair based on images generated by a Python script. Python-generated images are , and all others are resized with the longer side capped at 500 pixels.
@@ C.1 Model Access
This section provides details on model access and parameter settings (refer to Table 7). The model responses presented in this paper were collected between January 1 and February 28, 2025. We set , while all other parameters were kept at their default values.
@@
C.2 Detail of Evaluation
Prompt. Table 8 shows the prompts for evaluation. The one below is the prompt with CoT.
| Prompt |
| Question: {QUESTION} |
| Options: {OPTIONS} |
| Answer the question based on the most likely options. |
| Provide only the letter corresponding to your choice as the answer (e.g., ‘(a)’, ‘(b)’, ‘(c)’, ‘(d)’, ‘(e)’). |
| Question: {QUESTION} |
| Options: {OPTIONS} |
| Please think and answer the question based on the most likely options. |
Post-processing. We use InternVL2.5-38B to extract the selected options from the MLLMs response, and the corresponding extracting prompt is referenced from [16], as shown in Figure 10.
Human Evaluation. Two individuals with backgrounds in computer science, who were not involved in the project, independently participated by responding to the questions through a visualization interface. No compensation was provided for their participation.
C.3 Additional Results
@@
Table 9 presents the improvements introduced by the CoT prompt. It can be observed that, except for Gemini 2.0 Flash, which shows a positive gain on VisNumBench-Synthetic, the accuracy of the other models decreases. Tables 10 and 11 report the results of multimodal mathematical models and multimodal CoT models, respectively, corresponding to Figure 7. Table 12 presents the performance of models of varying sizes from the QwenVL and InternVL families, extending the results shown in Table 4. Table 13 summarizes the overall accuracy of various models on the VisNumBench benchmark, covering both synthetic and real subsets.
| Models | Angle | Length | Scale | Quantity | Depth | Area/Volume | Average |
| VisNumBench-Systentic | |||||||
| InternVL2.5-78B | 27.06 | 51.93 | 68.57 | 46.43 | 51.85 | 74.60 | 53.21 |
| Qwen2.5-VL-72B | 38.24 | 52.49 | 68.57 | 56.63 | 48.89 | 74.07 | 56.68 |
| Gemini 2.0 Flash | 34.12 | 55.80 | 85.00 | 58.16 | 55.56 | 74.60 | 60.14 |
| VisNumBench-Real | |||||||
| InternVL2.5-78B | 30.87 | 59.26 | 58.74 | 78.23 | 48.70 | 55.10 | 55.10 |
| Qwen2.5-VL-72B | 32.21 | 52.47 | 52.45 | 70.75 | 46.10 | 53.06 | 51.11 |
| Gemini 2.0 Flash | 33.56 | 49.38 | 72.03 | 80.95 | 41.56 | 59.86 | 55.88 |
@@
| Model | Synthetic (1,011) | Real (902) | Overall |
| Human | 95.33 | 97.33 | 96.27 |
| Gemini 2.0 Flash | 57.57 | 56.54 | 57.08 |
| InternVL2.5-78B | 56.18 | 56.54 | 56.35 |
| Qwen2.5-VL-72B | 58.46 | 53.33 | 56.04 |
| InternVL2.5-38B | 55.59 | 52.11 | 53.95 |
| LLaVA-Onevision-72B | 50.84 | 50.78 | 50.81 |
| Qwen2-VL-72B | 54.20 | 46.56 | 50.60 |
| LLaVA-v1.6-34B | 44.31 | 50.55 | 47.25 |
| Gemini 1.5 Pro | 44.02 | 48.67 | 46.21 |
| InternVL2-40B | 45.50 | 45.12 | 45.32 |
| Qwen2.5-VL-7B | 46.19 | 41.02 | 43.75 |
| Llama-VL-3_2-11B | 43.92 | 43.24 | 43.60 |
| Qwen2.5-VL-3B | 42.43 | 42.57 | 42.50 |
| Qwen2-VL-7B | 42.24 | 41.91 | 42.08 |
| Llama-3.2V-11B-cot | 45.50 | 38.36 | 42.13 |
| GPT-4o | 43.72 | 39.58 | 41.77 |
| InternVL2-8B-MPO | 40.65 | 39.91 | 40.30 |
| LLaVA-Onevision-7B | 39.96 | 40.58 | 40.25 |
| InternVL2.5-8B | 39.66 | 40.13 | 39.88 |
| InternVL2-8B | 39.56 | 39.58 | 39.57 |
| R1-Onevision-7B | 38.87 | 38.25 | 38.58 |
| LLaVA-v1.5-13B | 32.15 | 40.02 | 35.86 |
| Janus-Pro-7B | 37.69 | 34.26 | 36.07 |
| Phi-3.5-vision | 32.34 | 37.25 | 34.66 |
| Math-LLaVA-13B | 35.81 | 33.15 | 34.56 |
| Gemini 1.5 Flash | 33.33 | 33.70 | 33.50 |
| LLaVA-v1.5-7B | 29.38 | 28.49 | 28.96 |
| Qwen2-VL-2B | 31.85 | 24.94 | 28.59 |
| Random | 24.76 | 25.54 | 25.13 |
C.4 Error Analysis
We use the results of Gemini 2.0 Flash as a case study for error analysis. Based on these results, we randomly selected 10 erroneous instances for each attribute in each scenario, manually analyzing a total of 120 randomly sampled errors across all tasks. We categorize the errors into two types: (1) Errors arising from the model’s failure to accurately perceive the image (image perception errors); (2) Errors in which the model correctly interprets both the image and the question but fails to produce the correct numerical answer (numerical intuition errors). Our analysis reveals that of the errors are due to image perception issues, particularly in scale-related tasks, where the model struggles to identify the positions of pointers. The remaining are attributed to numerical intuition errors, which commonly involve difficulties with depth estimation, angle relationships, and quantity perception. These findings further substantiate that current models indeed lack robust numerical intuition.
Appendix D Example Data and Model Outputs